OpenStudy (anonymous):

will give medal Lena's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Lena $5.65 per pound, and type B coffee costs $4.35 per pound. This month, Lena made 167 pounds of the blend, for a total cost of $852.55 . How many pounds of type B coffee did she use? Lena's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Lena $5.65 per pound, and type B coffee costs $4.35 per pound. This month, Lena made 167pounds of the blend, for a total cost o$852.55. How many pounds of type B coffee did she use?

OpenStudy (mertsj):

x+y=167 5.65x+4.35y=852.55 Solve that system

OpenStudy (matlee):

x+y=167 5.65x+4.35y=852.55 Solve that system